# Facial Surgery Preview Based on the Orthognathic Treatment Prediction

### v1

@@ Abstract
Orthognathic surgery consultation is essential to help patients understand the changes to their facial appearance after surgery. However, current visualization methods are often inefficient and inaccurate due to limited pre- and post-treatment data and the complexity of the treatment. To overcome these challenges, this study aims to develop a fully automated pipeline that generates accurate and efficient 3D previews of postsurgical facial appearances for patients with orthognathic treatment without requiring additional medical images. The study introduces novel aesthetic losses, such as mouth-convexity and asymmetry losses, to improve the accuracy of facial surgery prediction. Additionally, it proposes a specialized parametric model for 3D reconstruction of the patient, medical-related losses to guide latent code prediction network optimization, and a data augmentation scheme to address insufficient data. The study additionally employs FLAME, a parametric model, to enhance the quality of facial appearance previews by extracting facial latent codes and establishing dense correspondences between pre- and post-surgery geometries. Quantitative comparisons showed the algorithm’s effectiveness, and qualitative results highlighted accurate facial contour and detail predictions. A user study confirmed that doctors and the public could not distinguish between machine learning predictions and actual postoperative results. This study aims to offer a practical, effective solution for orthognathic surgery consultations, benefiting doctors and patients.
@@ 1 Introduction
Orthognathic surgery corrects facial asymmetry and abnormalities, improving aesthetics, oral function, and psychosocial well-being. However, patients often feel anxious due to uncertainty about their post-surgical appearance, impacting communication with their doctors. Providing accurate previews of the expected outcome can help reduce pre-surgical anxiety, set realistic expectations, and increase patient satisfaction.
While Virtual Surgical Planning (VSP) tools (e.g. Dolphin [10], Morpheus 3D FaceMaker[4], and 3dMD Vultus [1]) provide clinical level high accuracy results for surgery planning use, they require Computed Tomography (CT) or Cone Beam Computed Tomography (CBCT) imaging data as additional inputs, which is always inaccessible for preview purpose. The introduction of CT/CBCT data has some practical limitations: It significantly increases the requirements of capturing data for users in search of functional or aesthetic improvements for jaws. On the other hand, due to the complexity of clinical simulations, this type of tools requires significant expertise and labor from physicians and cannot be fully automated. Therefore, the series of surgical planning tools cannot be used as preview tools for the general public.
Several machine learning-based solutions have been proposed for facial surgery preview purposes. These solutions utilize dense multilayer perceptrons (MLP) [25, 26, 3, 23, 20], or employ a conditional generative adversarial network [21, 13] to predict 3D coordinate changes. However, these solutions require X-ray or CT data and target positions of anatomical landmarks as additional input, which can only be specified by clinicians, making them not fully automated. Furthermore, the use of surgical planning parameters by medical professionals significantly increases the cost and complexity of these approaches. As a result, they may not be feasible for individuals seeking quick and intuitive surgical consultations. Currently, available clinician-oriented solutions only provide automated diagnosis and treatment simulations for clinical and research purposes, and there is no study focusing on automatic facial surgery preview specifically for patient visualization purposes.
A patient-oriented solution for visualizing surgical outcomes should be more accessible than a professional doctor-oriented one. In the early consultation stage, potential patients are not obligated to take CBCT or X-ray imaging, which would expose users to radiation. Instead, 3D facial reconstruction (e.g., 3dMD (3dMD Inc., Atlanta, GA, USA), Bellus3D (Bellus3D Inc., Los Gatos, CA, USA), and Ein-Scan3D (Shining 3D Technology, Hangzhou, China)) is much safer for post-treatment preview purpose. The solution should be fully automatic and efficient, providing prompt visualization to the user. An automatic facial surgery preview can serve as an effective educational and motivational tool for patients, in addition to serving as a visual aid in acceptance of the treatment.
@@
Our algorithm is designed to predict 3D post-surgery facial appearance based on preoperative facial scans of patients without the need for additional medical images (e.g., X-Ray or CBCT). To address the challenge posed by insufficient data, we develop a data augmentation method that substantially amplifies the available dataset for orthognathic treatment. To explicitly encode the assessment rule in orthognathic treatment, we introduce two medical losses, mouth-convexity loss and asymmetry loss, consistent with surgical definitions that help achieve the optimization goals of surgery. Moreover, our algorithm leverages the unique capabilities of FLAME [15, 30, 16], a parametric model, to extract facial features and build dense correspondences between pre- and post- facial geometries, which is crucial for texture transfer.
In summary, our contributions are: (1) We introduce two novel losses: mouth-convexity loss, and asymmetry loss, that align with the goal of orthognathic surgery, to improve prediction accuracy; (2) We propose an effective data augmentation scheme addressing the prevalent issue of insufficient data in medical practice; (3) We complete a fully automated pipeline that showcases the postoperative changes in facial appearance to patients in 3D.
2 Related Work
2.1 Virtual Surgery Planning for orthognathic surgery
Using machine learning techniques for surgical prediction has become a popular trend. Data-driven algorithms can be used to estimate the likelihood of surgical necessity and complexity, providing valuable information for surgeons to consider. The study of [11] shows that neural networks can determine the need for orthognathic surgery using only facial photographs. [27] presents a deep learning model to predict the difficulty in extracting teeth using radiographic images. However, these publications were not aiming at only predictions of probability, not predictions in the field of vision.
The researchers have then begun exploring the use of machine learning algorithms to predict facial appearance after surgery. These trials include linear regression [12], multilayer perceptrons [25, 26], and conditional GAN[21]. However, in [26] Rutger’s algorithm did not take into account the significant reduction in facial asymmetry before and after surgery, which may limit its ability to correct facial asymmetry and predict the appearance of asymmetric patients accurately, as Rutger acknowledged. In [25], Tanikawa’s network, trained on fewer than 100 patients, predicts postoperative displacement of dense point coordinates. However, the high dimensionality of the input and output layers (around 18,000) raises overfitting concerns. In addition, part of the network input is derived from CBCT data, which are difficult to obtain as previously mentioned. In [21], while cGAN has proven effective in predicting changes in facial appearance after orthodontic surgery, its applicability to predicting changes after orthognathic surgery may be limited. This is because orthognathic surgery involves greater complexity and a wider range of factors to consider compared to orthodontic surgery. [18] proposes a deep learning framework, FC-Net, the first deep learning-based method to predict postoperative facial changes in patients with skull-facial deformities. Its main purpose is for the doctors to use, as during the inference stage, the precise location of bone movement must be provided by the doctor as a reference in order to present the postoperative appearance.
2.2 Facial beautification
2D Beautification Techniques. While there are numerous recreational facial beautification applications, they are not suitable for presenting post-surgical results to patients due to their lack of medical foundation and surgical background. Most of these software are based on one-click beautification techniques that reflect popular aesthetics and do not allow fine manipulation by surgeons. Furthermore, the effects of commercial applications such as B612 [7], ModiFace [5], and YouCam [6] can exceed what can be achieved by orthognathic surgery. In contrast, our proposed algorithm is based on the displacement of vertices and can precisely beautify facial contour in a medically sound manner. 2D facial enhancement can also be achieved by using face-swap methods. FaceShifter [14], a two-stage framework that uses innovative encoding and refinement techniques to swap faces, is capable of facial beautification while preserving identity. FSGAN[19] presents a subject-agnostic face swapping and reenactment approach that does not require training on specific face pairs, featuring a novel iterative deep learning method for handling significant pose and expression variations. The subject-agnostic nature of the method makes it more accessible and user-friendly for people seeking medical consultation, since they do not need to provide training data for each pair of faces they want to swap or re-enact.
3D Beautification Techniques. 3D beautification can be achieved through 3D facial reconstruction, which has evolved from large-scale modeling (LSFM in [2]) to capturing details and customizing for specific individuals or groups, enhancing the overall visual appeal and aesthetics of the reconstructed face. In [17], the Structure-Aware Editable Morphable Model (SEMM) introduces a detailed structure representation, addressing the need to capture facial details in addition to overall geometry. SEMM provides detailed facial representation and intuitive wrinkle manipulation, enabling precise age-related feature control for realistic facial beautification. Detailed reconstructions are highly beneficial for predicting outcomes in medical aesthetics and plastic surgery, allowing more accurate and effective treatment planning. In terms of orthognathic surgery, Qiu proposed SCULPTOR in [22], a skeleton-consistent face generator that can predict facial changes consistent with skeletal variations, especially those resulting from dental surgery. Sculptor, as a refined 3DMM specifically designed for orthognathic surgery, could provide more effective predictions for virtual surgery planning if it were open source, benefiting practitioners and patients by improving surgical outcomes.
3 Methodology
@@
Figure 1: Net architecture for predicting the postoperative appearance from a captured 3D scan. During the training phase, the captured mesh and auto-annotated landmarks are first passed through the FLAME fitting procedure, where they are transformed into a compressed latent code. It is then passed through a predictor and is added into the output code difference. The FLAME reconstructing procedure helps to calculate well-designed losses using markers and mesh. With the help of medical, latent code, and geological types of loss, the parameter of the code difference predictor can be continuously updated. During the testing phase, the data do not flow through the dashed line. The predicted postoperative appearance is generated after FLAME reconstructing.
We present a simple and effective pipeline to predict 3D face models after surgery, as shown in Fig.1. With a reconstructed 3D facial model as input, we encode it as a latent code with the help of FLAME [15], a parametric facial model. Then we adopt a face predictor to predict the difference between the post- and pre- surgery faces under our well-tailored losses. Those losses not only consider the deviation between latent codes, but also explicitly formulate the facial-asymmetry and mouth-convexity rules of aesthetics in the orthognathic field. Those rules are computed on the predicted facial geometry directly and back-propagated to the network in a differentiable manner. Finally, owing to the mesh coherence in the FLAME model, as a better visualization, we transfer the texture from pretreatment 3D facial scans to our predicted face models to provide a fully textured post-treatment preview.
As we discussed earlier, to address the challenge of insufficient training data, we also propose a novel data augmentation approach to significantly increase the number of face pairs.
This study was conducted with the approval of the Institutional Review Board (IRB) under the joint agreement of the University of Hong Kong and Hospital Authority Hong Kong West Cluster. The IRB reference number for this study is UW 21-140, with approval granted on February 17, 2021. All procedures adhered to the Declaration of Helsinki and local regulatory requirements.
@@ 3.1 Prediction Network
The overall loss function used in our algorithm is composed of four components: the mouth convexity loss, the asymmetry loss, the latent code loss, and the geometry loss.
The main challenge in training the predictor comes from the imperfect training data. Our training data consist of real orthognathic surgery cases. However, due to some practical limitations or customized considerations, the post-treatment facial appearance may not be ideal from a clinical perspective, e.g. some of them are still asymmetric or protruding to a certain extent. Adopting a pure data-driven loss in supervision would yield the same artifacts as observed in the training data. To this end, we formulate two explicit clinical rules as our novel losses, the mouth-convexity loss and the asymmetry loss, to enhance the functional and aesthetic aspects.
@@ 3.2 Data Augmentation
A sufficient amount of diverse data is essential for updating machine learning model parameters [29], but in orthognathic surgery, data availability is limited by ethical issues and a small number of cases. Additionally, earlier data scarcity was due to the lack of advanced scanning devices, standardized procedures, and insufficient awareness among medical professionals to record 3D facial data. As a result, data augmentation is vital for predicting orthognathic 3D facial outcomes.
Traditional $3$D data augmentation methods like translation, rotation, and scaling may not be effective in our framework without generating meaningful latent codes. Therefore, we propose creating new $3$D facial meshes by combining a patient’s lower face with a randomized upper face. The challenge is the difficulty in seamlessly splicing the two due to differences in facial proportions and features. To address this, we design a novel data augmentation algorithm that utilizes FLAME to fit the lower part of postoperative face. Since the fitting process is under-constrained, it enables the generation of multiple outcomes while preserving the same lower face. We perform data augmentation on pairs of pre- and post-surgical scans. For the postoperative side, the data is seamlessly augmented with the synthetic upper face. For the preoperative side, since the surgery only alter the lower face, we stitch the lower part of the preoperative scan with the synthetic upper part to create a seamless model. We also include a data cleaning step to remove outliers caused by the fitting process. By calculating the fitting error of key facial landmarks and setting a threshold, only meshes within the threshold are retained, ensuring the accuracy and reliability of the synthetic meshes.
@@ 3.3 Integrating FLAME as an encoder-decoder
FLAME, as an encoder, offers better fit accuracy than 3DMM, a linear model based on PCA, due to its incorporation of the jaw joint in its skeleton tree, which captures jaw movements from orthognathic surgeries. As a decoder, FLAME reconstructs facial geometry from the predicted latent code. Geometric information is used to calculate loss during training and optimize the latent code through backpropagation. During inference, the full-face mesh output helps render a textured 3D face model.
3.4 Post processing: scan deformation
We aim to provide people who come for surgical consultations with a textured 3D mesh that can be viewed from multiple angles, giving them a comprehensive and visual understanding of their postoperative appearance. To achieve this, we deform the captured facial scan so that it conforms to the predicted FLAME mesh. We morph the dense grid by deforming the sparse grid. We first build the correspondences between the dense and sparse grids. For each vertex in the dense mesh, we find the closest point in the sparse mesh as its correspondence, whose barycentric coordinate is $(\lambda_{1},\lambda_{2},\lambda_{3})$ and the morphing displacement of the triangular vertex belonging to it is $d_{1}$, $d_{2}$, and $d_{3}$. The morphing process is based on the barycentric coordinates, i.e. $\lambda_{1}d_{1}+\lambda_{2}d_{2}+\lambda_{3}d_{3}$.
4 Experiments
4.1 Data Collection and Pre-processing
Data Collection. Using the 3dMD facial scan acquisition system, $163$ pairs of pre- and post-operative 3D facial scans of orthognathic surgery were collected from the University of Hong Kong School of Dentistry. Postoperative scans were taken at least $3$ months after the surgical procedure and were used to ensure that the swelling of the patient had subsided and could truly reflect the expected appearance of the patient after surgery. The ratio of male to female patients is $59:104$, with the majority of Asian patients.
Data Cleaning. Our analysis focuses solely on the face, excluding non-facial elements like hair and hats. This ensures that our algorithm can concentrate on key facial areas, such as the chin and nose, while also improving FLAME’s fitting accuracy, as it excludes hair and accessories. To remove non-facial parts, we render the scans from three viewpoints: frontal and both side views with a 45-degree rotation. We then use BiSeNet [28], a bilateral segmentation network, to parse the facial regions from these images, retaining only the relevant areas for analysis.

### v2

@@ Abstract
Background and Objective: Orthognathic surgery consultations are essential for helping patients understand how their facial appearance may change after surgery. However, current visualization methods are often inefficient and inaccurate due to limited pre- and post-treatment data and the complexity of the treatment. This study aims to develop a fully automated pipeline for generating accurate and efficient 3D previews of postsurgical facial appearances without requiring additional medical images.
Methods: The proposed method incorporates novel aesthetic criteria, such as mouth-convexity and asymmetry, to improve prediction accuracy. To address data limitations, a robust data augmentation scheme is implemented. Performance is evaluated against state-of-the-art methods using Chamfer distance and Hausdorff distance metrics. Additionally, a user study involving medical professionals and engineers was conducted to evaluate the effectiveness of the predicted models. Participants performed blinded comparisons of machine learning-generated faces and real surgical outcomes, with McNemar’s test used to analyze the robustness of their differentiation.
Results: Quantitative evaluations showed high prediction accuracy for our method, with a Hausdorff Distance of 9.00 millimeters and Chamfer Distance of 2.50 millimeters, outperforming the state of the art. Even without additional synthesized data, our method achieved competitive results (Hausdorff Distance: 9.43 millimeters, Chamfer Distance: 2.94 millimeters). Qualitative results demonstrated accurate facial predictions. The analysis revealed slightly higher sensitivity (54.20% compared to 53.30%) and precision (50.20% compared to 49.40%) for engineers compared to medical professionals, though both groups had low specificity, approximately 46%. Statistical tests showed no significant difference in distinguishing Machine Learning-Generated faces from Real Surgical Outcomes, with p-values of 0.567 and 0.256, respectively. Ablation tests demonstrated the contribution of our loss functions and data augmentation in enhancing prediction accuracy.
Conclusions: This study provides a practical and effective solution for orthognathic surgery consultations, benefiting both doctors and patients by improving the efficiency and accuracy of 3D postsurgical facial appearance previews. The proposed method has the potential for practical application in pre-surgical visualization and aiding in decision-making.
@@
{graphicalabstract}[Uncaptioned image]
{highlights}
We propose an automated method for 3D post-surgical prediction using only multi view images.
We integrate mouth convexity and asymmetry criteria to enhance orthognathic planning.
We generate synthetic data from real cases to enhance training, improving robustness.
@@ 1 Introduction
Orthognathic surgery addresses facial asymmetry and abnormalities, significantly improving aesthetics, oral function, and psychosocial well-being. Despite these benefits, uncertainty about the appearance of the postoperative period often leads to anxiety among patients, affecting their decision-making process and communication with doctors. Visualizing expected surgical outcomes has emerged as a crucial step in mitigating presurgical anxiety, setting realistic expectations, and improving overall patient satisfaction, particularly when CBCT data is not available.
Data-driven approaches have shown promise in assessing surgical necessity and complexity. For example, neural networks have been used to predict the need for orthognathic surgery from facial photographs [1] and to estimate the difficulty of tooth extraction from radiographic images [2]. However, these studies focus primarily on probability predictions rather than visual outcomes.
To address the need for predictive visualizations, researchers have begun exploring the use of machine learning algorithms to predict facial appearance after surgery [3]. These algorithms employ various advanced methods, including dense multilayer perceptrons (MLP)[4, 5, 6, 7, 8, 9, 10], conditional generative adversarial networks (cGAN)[11, 12], transformers [13], and convolutional neural networks (CNN) [14]. These deep learning models excel at capturing spatial hierarchies in images, enabling refined feature extraction and accurate prediction of 3D coordinate changes. Despite these advancements, existing methods remain clinician-oriented, heavily reliant on X-ray or CT data and anatomical landmarks provided by medical professionals [15]. This reliance limits full automation and increases costs, making these approaches inaccessible for patients seeking quick and intuitive consultations. Moreover, several limitations persist. For example, the model in [6] overlooks key reductions in facial asymmetry after surgery, potentially compromising accuracy for asymmetric patients. Similarly, Tanikawa’s network [5], trained on fewer than 100 patients, faces overfitting issues due to high-dimensional input and output layers (18,000), while requiring CBCT data, which are challenging to obtain. Other approaches, such as Park’s cGAN [11], are effective for orthodontic surgery but lack robustness for orthognathic procedures due to their complexity. Finally, methods like FC-Net [16] require clinicians to provide precise bone movement data during inference, limiting their usability for direct patient engagement.
To accurately predict 3D post-surgery facial appearance, it is essential to use parametric 3D facial reconstruction methods that capture fine articulation, especially the jaw joint, which is critical for both aesthetic and functional outcomes in orthognathic surgery. Existing methods like Large Scale Facial Model (LSFM) [17] and Structure-Aware Editable Morphable Model (SEMM) [18] fail to model dynamic jaw movement, treating the face as a morphable structure where all features, including the jaw, are considered as a whole. SCULPTOR [19], an articulated model, can capture skull-face joint distribution but is difficult to integrate due to limited data and incomplete code. To overcome this, we integrate FLAME [20, 21, 22], a model with nonlinear jaw articulation, enabling vivid facial reconstruction and efficient encoding for accurate post-surgical predictions.
@@
Our algorithm is designed to predict 3D post-surgery facial appearance based on preoperative facial scans of patients without the need for additional medical images (e.g., X-Ray or CBCT). To improve the robustness of the model, we develop a data augmentation method that substantially amplifies the available dataset for orthognathic treatment. To explicitly encode the assessment rules in orthognathic treatment, we introduce two medical losses, mouth-convexity loss and asymmetry loss, consistent with surgical definitions that help achieve the optimization goals of surgery.
In summary, our contributions are: (1) We integrate two criteria—minimizing mouth-convexity and minimizing asymmetry—into the machine learning procedure, which align with the goals of orthognathic surgery, to enhance prediction accuracy; (2) we devise an augmentation method to expand the available dataset, achieving more robust predictions; (3) We complete a fully automated pipeline that showcases the postoperative changes in facial appearance to patients in 3D.
2 Methods
@@
Figure 1: Net architecture for predicting the postoperative appearance from a captured 3D scan. During the training phase, the captured mesh and auto-annotated landmarks are first passed through the FLAME fitting procedure, where they are transformed into a compressed latent code. It then passes through a predictor and has the code difference added to it. The FLAME reconstructing procedure helps to calculate well-designed losses using markers and mesh. With the help of medical, latent code, and geological types of loss, the parameter of the code difference predictor can be continuously updated. During the testing phase, the data do not flow through the dashed line. The predicted postoperative appearance is generated through reconstruction using FLAME.
We present a simple and effective pipeline to predict 3D face models after surgery, as shown in Fig.1. The reconstructed 3D facial model is initially encoded into a latent code using FLAME [20], a parametric facial model. A face predictor then estimates the difference between post- and pre-surgery facial features, guided by custom-tailored loss functions that specifically incorporate facial asymmetry and mouth-convexity rules relevant to orthognathic surgery (explained in Section 2.1). These rules are directly applied to the predicted facial geometry and back-propagated through the network in a differentiable manner (discussed in Section 2.3). To provide a fully textured post-treatment preview, we leverage the mesh coherence of the FLAME model, transferring the texture from pre-treatment 3D facial scans to the predicted face models (explained in Section 2.4). To address the challenge of limited training data, we introduce a novel data augmentation technique that significantly increases the number of face pairs, enhancing the robustness of the model (described in Section 2.2). The effectiveness of the predicted models is evaluated through a user study involving both medical professionals and engineers, with the robustness of the predictions assessed using McNemar’s test (outlined in Section 2.5).
@@ 2.1 Prediction network
The overall loss function used in our algorithm is composed of four components: the mouth convexity loss, the asymmetry loss, the latent code loss, and the geometry loss. The main challenge in training the predictor comes from the imperfect training data. Our training data consist of real orthognathic surgery cases. However, due to some practical limitations or customized considerations, the post-treatment facial appearance may not be ideal from a clinical perspective, e.g. some of them are still asymmetric or protruding to a certain extent. Adopting a pure data-driven loss in supervision would yield the same artifacts as observed in the training data. To this end, we formulate two explicit clinical rules as our novel losses, the mouth-convexity loss and the asymmetry loss, to enhance the functional and aesthetic aspects.
@@ 2.2 Data augmentation
A sufficient amount of diverse data is crucial for enhancing the robustness of machine learning models [25]. In orthognathic surgery, data augmentation offers significant advantages by improving the algorithm’s ability to generalize across a wide range of cases. By simulating diverse populations, data augmentation enhances the model’s predictive accuracy and ensures reliable 3D facial outcome predictions.
A common assumption in orthognathic surgery is that there is a horizontal plane dividing the unchanged upper part of the face (above the plane) from the key modified region (below the plane), where the surgical alterations occur. The goal of our data augmentation method is to generate synthetic data where the upper face remains intact, preserving its natural characteristics above the horizontal plane, while the lower part, which is affected by the surgery, is modified according to the surgical requirements. Traditional 3D data augmentation techniques, such as translation, rotation, and scaling, may not be effective in our framework, as they do not generate meaningful geometry that meets the requirements for preserving the natural structure and characteristics of the orthognathic surgery.
Figure 3: An illustration of synthetic data generation process. The process involves three stages: face segmentation, conditionally generating a synthetic face $F_{gen}$, and stitching the upper and lower regions to create plausible pre- and post-surgery pairs for training.
To create the synthetic data, we first identify the horizontal plane $P_{s}$ that separates the unaltered upper face from the altered lower region. As illustrated in Figure 3, for the pre-surgery scan, we obtain $F_{pre}^{u}$ (upper part) and $F_{pre}^{l}$ (lower part). For the post-surgery scan, we obtain $F_{post}^{u}$ (upper part) and $F_{post}^{l}$ (lower part). Note that $F_{pre}^{u}$ is the same as $F_{post}^{u}$. Next, we conditionally generate a synthetic face $F_{gen}$ based on the lower part of the postoperative scan $F_{post}^{l}$ with a random variable . Then, we use the horizontal plane $P_{s}$ to separate the upper part of $F_{gen}$, denoted as $F_{gen}^{u}$. Finally, we create the pair of generated faces $\{\tilde{F}_{pre},\tilde{F}_{post}\}$ by stitching $F_{gen}^{u}$ with both $F_{pre}^{l}$ and $F_{post}^{l}$, respectively. We use a random variant $\xi$ to alter the latent code of the upper face, and then we use the decoder to reconstruct the facial geometry. We randomize the $\xi$ to synthesize multiple plausible pairs of $\{\tilde{F}_{pre},\tilde{F}_{post}\}$ to train the prediction network. This approach ensures that the synthetic data reflects the surgical changes below the horizontal plane while keeping the upper part of the face unchanged.
In summary, our goal is to generate paired facial scans that reflect the pre- and post-surgery appearance. We modify the upper part of the face while directly copying the lower part from the dataset, following the paired changes between the pre- and post-surgery scans. As a result, we obtain a pair of faces where the upper part remains consistent, and the lower part exhibits the surgical changes.
Additionally, we include a data cleaning step to remove outliers introduced by the fitting process. By calculating the fitting error at key facial landmarks and applying a threshold, we ensure that only meshes with an acceptable error are retained, thereby guaranteeing the accuracy and reliability of the synthetic meshes. To evaluate the impact of data augmentation, we conducted an ablation study, with the results presented in the last row of Table 6 and the last row of Figure 5
@@ 2.3 Integrating FLAME as an encoder-decoder
2.4 Visualization of prediction as a textured mesh
We aim to provide individuals seeking surgical consultations with a textured 3D mesh that can be viewed from multiple angles, offering a comprehensive and intuitive visualization of their postoperative appearance.
To achieve this, we deform the textured scan to align with the geometry of the predicted parametric model. Specifically, using barycentric coordinates, we deform the vertices to match the geometry of the predicted model. For each vertex in the scanned mesh, the closest point in the parametric model is identified as its correspondence, and the displacement for each vertex is computed based on the barycentric coordinates $(\lambda_{1},\lambda_{2},\lambda_{3})$ and their predicted displacements $d_{1}$, $d_{2}$, and $d_{3}$. The texture is then morphed using the formula $\lambda_{1}d_{1}+\lambda_{2}d_{2}+\lambda_{3}d_{3}$ to approach the geometry of the predicted parametric model.
2.5 User study
The aim of this study was to assess the participants’ ability to differentiate between machine learning-generated faces (MLG) and real surgical outcomes (RSO). Participants in this study consisted of five medical professionals (comprising two surgeons and three orthodontists) with a minimum of 5 years of experience in their respective fields. Additionally, 15 engineers (including twelve males and three females) were included in the study, each possessing a minimum of 3 years of relevant experience in engineering and technical disciplines. The selection criteria for medical professionals were based on their expertise in surgical procedures and orthodontic treatments, ensuring a minimum threshold of 5 years of clinical experience. Engineers were chosen for their technical background and familiarity with machine learning concepts, with a minimum of 3 years of professional experience in their respective fields. Preoperative and post-surgery 3D facial data were collected for each patient, standardized using FLAME parameters, and textures were applied consistently across all data. To facilitate participants’ observation of facial symmetry and feature positioning, both frontal and side views of the faces were provided, along with a 180-degree rotating animation from left to right. The preoperative images were clearly labelled, while the MLG and RSO were presented to the participants in a blinded manner. Each participant was simultaneously shown two images (labelled as A and B) and asked to distinguish between MLG and RSO without knowing the group to which each image belonged. A total of 30 randomly selected images (including A and B) were presented to each participant. Participants were required to digitally record their responses. The responses provided by the participants were then compared with the correct answers separately for engineers and medical professionals, and specificity, sensitivity, and precision values were calculated for each group. To further analyze the participants’ ability to differentiate between MLG and RSO, a McNemar’s test was conducted for each group.
3 Results
3.1 Datasets
A total of 163 pairs of pre- and post-operative 3D facial scans were collected using the 3dMD facial scan acquisition system from patients undergoing orthognathic surgery at the University of Hong Kong School of Dentistry. Static three-dimensional (3D) images of each participant were taken using a 3dMDface System (3dMD Inc., Atlanta, GA, USA) by professional photographers. The accuracy of the system had been previously published and was reported to be lower than 0.2 mm root mean square (RMS) [26]. The system was calibrated according to the manufacturer’s instructions before each image capture. During scanning, patients were relaxed and instructed to look straight into a mirror at their own eyes, ensuring the capture of their forehead, chin, and ears. Immediately prior to image capture, participants were seated 100 cm away from the system, looking forward with the Frankfort plane parallel to the floor, and any glasses and jewelry were removed. The camera system captured six 2D images—four black-and-white pictures to depict facial structures and spatial relationships, and two colored images to project texture information onto the mesh framework. The scan process took 1.5 milliseconds and the 3D facial surfaces were exported as Wavefront OBJ files for further processing. Postoperative scans were recorded at least three months after the surgery to ensure that facial swelling had subsided, providing an accurate reflection of the patient’s post-recovery appearance. The male-to-female ratio was 59:104, with the majority of patients being of Asian descent.
3.2 Implementation details
3.2.1 Exclusion of non-Facial elements
Data cleaning is a crucial step in our prediction algorithm, as it ensures that the focus remains solely on the relevant facial regions by deliberately excluding non-facial elements such as hair and disposable surgical caps. This process is essential to eliminate extraneous data that could introduce noise, allowing the algorithm to concentrate on critical facial features like the chin and nose. By removing these non-facial elements, we not only enhance the fitting accuracy but also improve the overall performance of the model. The careful segmentation and exclusion of non-facial data ensure that the algorithm’s analysis is based on high-quality, relevant information, leading to more precise and reliable predictions.
To isolate the facial regions, we render the 3D scans from three distinct viewpoints: frontal and both side views with a 45-degree rotation. We then apply BiSeNet [27], a bilateral segmentation network, to segment the facial regions from these images, retaining only the pertinent areas for subsequent analysis. Once the 2D segmentation maps are obtained, we reproject them back into the 3D space to ensure that the segmented regions correspond accurately to the facial areas in the original 3D scans. The non-facial elements, identified in the segmentation, are then removed from the 3D scans, leaving only the relevant facial regions for further analysis.
3.2.2 Landmarks annotation